Derive

the MU function from the following TU function:

๐‘‡๐‘ˆ = 200๐‘„ โˆ’ 25๐‘„2 + ๐‘„3

From this MU function, draw up a table up to the level of Q where MU becomes negative. i.e (Q = 1, 2, etc). Graph these figures. (10 marks)



1
Expert's answer
2021-04-23T10:42:23-0400

"TU=200Q-25Q^{2}+Q^{3}"

"MU=\frac{dTU}{dQ}"

"=\frac{d}{dQ}(200Q-25Q^{2}+Q^{3})" ย ย 

ย "=200-50Q+3Q^{2}"

Hence,ย 

"MU=200-50Q+3Q^{2}"

Table

Q Mu

1 153

2 112

3 77

4 48

5 25

6 8

7 -3

8 -8

Graph

The quantity is on the horizontal axis and the marginal utility on the vertical axis .


The marginal utility is convex to origin and downward sloping curve indicating that utility decreases at an increasing rate with each additional unit of consumption.ย 

ย